find the equation of the line through the point (1,-6)and which is parallel to the line 3x-y+4=0
Answer by LinnW(1048)

Step one is solve 3x-y+4=0 for y
add y to each side
3x+4=y
In the template y = mx + b, m is the slope
so our slope = 3
We want a line with slope 3 going through (1,-6)
So we will substitute 3 for m, 1 for x and -6 for y in y = mx + b
-6 = 3(1) + b
-6 = 3 + b
add -3 to each side
-9 = b
So the equation parallel to 3x-y+4=0 is y=3x-9